OpenStudy (anonymous):

how to find the integral of 1/(x^(3)sqrt(x^2-1))

sam (.sam.):

What method you have in mind?

sam (.sam.):

\[\int\limits \frac{1}{x^3 \sqrt{x^2-1}} \, dx\]

OpenStudy (anonymous):

First impression: trig sub. I haven't worked it out for myself, but that's what I would try.

OpenStudy (anonymous):

Partial fractions?

OpenStudy (anonymous):

Oh no, I suppose we do use the u-substitution

sam (.sam.):

I dont think you can u-sub, trig sub will be the easiest way

OpenStudy (anonymous):

oh and how would you do that?

sam (.sam.):

Let \[x=asec(\theta)\]

OpenStudy (anonymous):

oh Okay. That makes sense. I'll try to work it out. Thank you

sam (.sam.):

When you worked all those trigs you'll end up \[\int\limits \cos^2 \theta d \theta\] Then use \(\cos(2\theta)=1-2\cos^2(\theta)\) and simplify

OpenStudy (anonymous):

This is what I should get \[\frac{ \sqrt{(x ^{2}-1)} }{ 2x ^{2} }\]-\[\frac{ 1 }{ 2 }\arctan(\frac{ 1 }{ \sqrt{x ^{2}-1} })\]